# LaDiR: Latent Diffusion Enhances LLMs for Text Reasoning

### v3

@@ Abstract
Large Language Models (LLMs) demonstrate their reasoning ability through chain-of-thought (CoT) generation. However, LLM’s autoregressive decoding may limit the ability to revisit and refine earlier tokens in a holistic manner, which can also lead to inefficient exploration for diverse solutions. In this paper, we propose LaDiR (Latent Diffusion Reasoner), a novel reasoning framework that unifies the expressiveness of continuous latent representation with the iterative refinement capabilities of latent diffusion models for an existing LLM. We first construct a structured latent reasoning space using a Variational Autoencoder (VAE) that encodes text reasoning steps into blocks of thought tokens, preserving semantic information and interpretability while offering compact but expressive representations. Subsequently, we utilize a latent diffusion model that learns to denoise a block of latent thought tokens with a blockwise bidirectional attention mask, enabling longer horizon and iterative refinement with adaptive test-time compute. This design allows efficient parallel generation of diverse reasoning trajectories, allowing the model to plan and revise the reasoning process holistically. We conduct evaluations on a suite of mathematical reasoning and planning benchmarks. Empirical results show that LaDiR consistently improves accuracy, diversity, and interpretability over existing autoregressive, diffusion-based, and latent reasoning methods, revealing a new paradigm for text reasoning with latent diffusion. [^1] Code is available at https://github.com/mk322/LaDiR.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable reasoning abilities through extensive pretraining on human languages, yet the inherent limitations of the autoregressive (AR) paradigm are becoming increasingly difficult to overlook (Zhou et al., 2024b; Bachmann & Nagarajan, 2025). As shown in Fig. 1 (top left), their sequential nature prevents revising earlier tokens, making self-refinement inefficient and difficult  (Chen et al., 2024; Huang et al., 2024). Moreover, AR models generate a linear chain of thought (CoT) (Dziri et al., 2023; Wei et al., 2023), which limits reasoning diversity and restricts exploration of multiple valid solutions (Naik et al., 2024; Yu et al., 2024).
Diffusion models (Ho et al., 2020), originally introduced for generation in continuous domains like images, have recently gained attention in text generation for their ability to maintain global coherence and enable iterative refinement (Ye et al., 2024b; Nie et al., 2025; Lou et al., 2023; Yu et al., 2025c; Weligalle, 2025; Sahoo et al., 2024; Gulrajani & Hashimoto, 2023; Lovelace et al., 2024). Existing works largely emphasize the parallelization properties of diffusion models (Israel et al., 2025; Nie et al., 2025; Weligalle, 2025). Arguably, a more important direction is to ask: How can these approaches enhance the reasoning capabilities of LLMs? We focus on one particularly promising capability: the ability to self-correct and refine reasoning chains at semantic levels in latent space. As shown in Fig. 1 (top right), such self-refinement cannot be achieved by discrete diffusion language models that merely transit into masked tokens.
@@
To address this limitation, we introduce LaDiR (Latent Diffusion Reasoner), a flexible reasoning framework that encodes high-level semantic representations of reasoning steps into continuous latent tokens via a Variational Autoencoder (VAE) as latent thought tokens, and trains a latent diffusion model over them to perform reasoning. This bridges the gap between surface-level token refinement and deeper semantic reasoning. After the reasoning process, the model generates final answer tokens conditioned on the generated latent thought tokens.
Our proposed paradigm establishes a new reasoning framework as a post-training method, bringing several distinctive advantages. First, the iterative refinement ability of diffusion enables a better trade-off between accuracy and test-time compute, as additional denoising steps can be flexibly allocated to improve performance. Second, unlike autoregressive models, our approach leverages diffusion models for parallel exploration over blocks of latent tokens and applies repulsion to push examples in the batch apart during inference, enabling the generation of multiple diverse reasoning trajectories. Finally, leveraging a VAE-based latent space enhances interpretability over continuous diffusion models, making the reasoning process more transparent and readable.
Experimentally, we demonstrate that diffusion-based latent reasoning is not only more accurate but also qualitatively different from prior approaches. On math reasoning benchmarks, including GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), where Coconut (Hao et al., 2024) fails to surpass AR CoT supervised finetuning (SFT), LaDiR consistently outperforms it on average across 7 benchmarks with the LLaMA 3.1 8B model (Dubey et al., 2024). This suggests that modeling reasoning at the semantic level, rather than at the token level, may lead to more faithful intermediate steps that accumulate into stronger final answers. Moreover, on the Countdown planning task, LaDiR shows over 30% absolute improvement in both Pass@1 and Pass@100, indicating that latent thought tokens potentially enhance global planning ability, while parallel diversity exploration enables the model to generate diverse reasoning paths. Together, these findings suggest that diffusion-based latent reasoning provides a principled way to balance accuracy and diversity—key ingredients for advancing beyond sequential autoregressive reasoning.
@@ 2 Preliminaries
This section introduces key concepts and notations in VAE and latent diffusion models (Rombach et al., 2022). Detailed formulations and background information are provided in Appendix B.
@@ 2.1 Variational Autoencoder
A Variational Autoencoder (VAE) (Kingma & Welling, 2013) learns a latent representation of data by balancing reconstruction accuracy and prior regularization. We adopt the $\beta$-VAE (Higgins et al., 2017), where a scaling factor $\beta$ controls this trade-off:
@@ 2.2 Latent Diffusion and Flow Matching
Latent diffusion models (Ho et al., 2020; Rombach et al., 2022) generate data by denoising latent variables from Gaussian noise in the latent space of a VAE, which preserves high-level semantic structure. Diffusion can also be viewed as a continuous-time generative flow trained via flow matching (Lipman et al., 2022), which we adopt as our primary framework for its superior performance (see Appendix C). The training and inference processes are as follows:
@@ 2.3 Block Diffusion
To support flexible and variable-length sequence generation, we employ a block diffusion scheme (Arriola et al., 2025) that integrates autoregressive modeling with diffusion. Instead of applying diffusion to individual latent tokens or full sequence, the sequence is divided into contiguous blocks, and diffusion is performed at the block level. This hybrid design retains the open-ended generation of autoregressive models while introducing global coherence within each block. See Appendix B.4 for details.
@@ 3.3 Reasoning Model Training
While $f_{\psi}$ learns to predict latent reasoning trajectories, to avoid explicitly decoding these steps through the VAE decoder for efficiency during inference, we use the same model to autoregressively predict answer text tokens conditioned on the latent reasoning blocks. To this end, given the question $q$, the reasoning blocks $\mathbf{Z}^{(\leq B)}$, and the past answer tokens $y_{<w}$, the model predicts the next answer token $y_{w}$ with distribution $p_{\psi}(y_{w}\mid q,\mathbf{Z}^{(\leq B)},y_{<w}).$ The training objective for those answer tokens is the cross-entropy loss:
@@
To explicitly control the number of latent blocks, we introduce a special binary classification head on top of the LLM transformer backbone. It predicts whether the next block begins with a <SOA> (start-of-answer) or <BOT> (begin-of-thought) token whenever an <EOT> (end-of-thought) token is generated. Formally, let $\tau$ index positions of <EOT> tokens in the output. For each $\tau$, the model produces a distribution $p_{\psi}(s_{\tau}\mid q,\mathbf{Z}^{(\leq B)},y_{\leq\tau}),s_{\tau}\in\{\texttt{<SOA>},\texttt{<BOT>}\},$ and we minimize the corresponding classification loss:
$\mathcal{L}_{\mathrm{Spec}}=-\sum_{\tau\in\mathcal{T}_{\text{EOT}}}\log p_{\psi}(s_{\tau}\mid q,\mathbf{Z}^{(\leq B)},y_{\leq\tau}).$ (4)
@@ 3.3.2 Stage 2: Rollout training
After Stage 1, there is a mismatch between training and inference. During inference, the model must be conditioned on previous self-generated latents without access to oracle latents, suffering from error accumulation issue. To address this issue, Stage 2 adopts an rollout training. We keep the same number of blocks $B$ as in the ground truth, but instead of conditioning on oracle latents, the model generates its own latents $\tilde{\mathbf{Z}}^{(1:B)}$ from random noise using a fewer denoising steps (i.e., $50\rightarrow 10$, following FlowGRPO (Liu et al., 2025)). We keep the gradients on $\tilde{\mathbf{Z}}^{(1:B)}$ during denoising, allowing answer supervision to backpropagate through the trajectory and directly shape latent predictions. To avoid latent collapse as in Coconut w/o curriculum learning (Hao et al., 2024), we keep the flow matching loss. Therefore, the training objective is same as Eq. 5.
@@ 3.4 Reasoning Model Inference
    Finally, the diversity-guided prediction combines the base model output with the repulsion gradient, in a form analogous to classifier-free guidance (Ho & Salimans, 2022): $\hat{z}_{t-1}\;=\;f_{\psi}(\mathbf{x}_{t},t,x)\;+\;\gamma_{t}\,\mathbf{F}(z)$, where $f_{\psi}(\mathbf{x}_{t},t,x)$ is the model’s prediction at step $t$.
@@ 4.1 Mathematical Reasoning
We fine-tune pretrained LLMs on the DART-MATH dataset (Tong et al., 2024b), a large-scale dataset synthesized to enhance mathematical reasoning. For evaluation, we adopt two in-domain benchmarks, Math (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021), and five out-of-domain benchmarks to assess generalization: College-Math (Tang et al., 2024b), DeepMind-Math (Saxton et al., 2019), OlympiaBench-Math (He et al., 2024), TheoremQA (Chen et al., 2023), and Fresh-Gaokao-Math-2023 (Tang et al., 2024b). Detailed dataset descriptions are provided in Appendix E.
@@
We compare our approach against a diverse set of reasoning methods spanning both autoregressive and diffusion-based paradigms. For autoregressive models, we include Sol-Only, trained only on question–solution pairs without intermediate reasoning steps, and CoT, trained with full chain-of-thought supervision. We further evaluate several latent reasoning methods: Implicit CoT (iCoT) (Deng et al., 2023), which gradually removes explicit CoT tokens through curriculum learning; Pause Token (Goyal et al., 2023), which introduces a learnable pause token to provide additional computation before answering; Coconut (Hao et al., 2024), which leverages hidden states as latent reasoning tokens through curriculum learning; and Discrete Latent (Su et al., 2025), which compresses a span of text (e.g., 16 tokens) into discrete latent codes via VQ-VAE for reasoning. For diffusion-based language models, we compare with the open-sourced LLaDA 8B (Nie et al., 2025), evaluated both with and without SFT. We utilize LLaMA-3.1 8B (Dubey et al., 2024) as the backbone model for our framework as well as for the AR baselines for fair comparison.
@@
Diffusion Language Models - LLaDA 8B
@@ Results
As shown in Table 1, our method achieves the strongest overall performance on average, improving the average pass@1 accuracy by 2% over the best prior latent approach. Compared to text-based CoT baselines, our latent reasoning consistently yields more robust solutions, particularly on harder benchmarks such as DM-Math and College-level datasets, where direct text reasoning often struggles with long-horizon consistency. This suggests that reasoning in a latent space at semantic level learns more abstract reasoning patterns. Compared to prior latent approaches (e.g., Coconut), the latent diffusion objective provides a more principled objective for modeling continuous trajectories, leading to stronger generalization to out-of-domain settings such as TheoremQA. Also, incorporating the stage 2 training notably improves performance across all benchmarks, showing its effectiveness in mitigating error accumulation. Taken together, these results indicate that LaDiR combines the interpretability benefits of CoT-style reasoning (see Appendix D) and expressiveness of continuous latent space, producing generalizable reasoning traces.

### v4

@@ Abstract
Large Language Models (LLMs) demonstrate their reasoning ability through chain-of-thought (CoT) generation. However, LLM’s autoregressive decoding may limit the ability to revisit and refine earlier tokens in a holistic manner, which can also lead to inefficient exploration for diverse solutions. In this paper, we propose LaDiR (Latent Diffusion Reasoner), a novel reasoning framework that unifies the expressiveness of continuous latent representation with the iterative refinement capabilities of latent diffusion models for an existing LLM. We first construct a structured latent reasoning space using a Variational Autoencoder (VAE) that encodes text reasoning steps into blocks of thought tokens, preserving semantic information and interpretability while offering compact but expressive representations. Subsequently, we utilize a latent diffusion model that learns to denoise a block of latent thought tokens with a blockwise bidirectional attention mask, enabling longer horizon and iterative refinement with adaptive test-time compute. This design, combined with explicit diversity guidance during diffusion inference, enables the generation of multiple diverse reasoning trajectories that explore distinct regions of the latent space, rather than producing repetitive solutions as often occurs in standard autoregressive sampling. We conduct evaluations on a suite of mathematical reasoning and planning benchmarks. Empirical results show that LaDiR consistently improves accuracy, diversity, and interpretability over existing autoregressive, diffusion-based, and latent reasoning methods, revealing a new paradigm for text reasoning with latent diffusion.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable reasoning abilities through extensive pretraining on human languages, yet the inherent limitations of the autoregressive (AR) paradigm are becoming increasingly difficult to overlook (zhou2024surveyefficientinferencelarge; bachmann2025pitfallsnexttokenprediction). As shown in Fig. 1 (top left), their sequential nature prevents revising earlier tokens, making self-refinement inefficient and difficult  (chen2024not; huang2024largelanguagemodelsselfcorrect). Moreover, AR models with discrete CoT generate a linear chain of thought (CoT) (dziri2023faithfatelimitstransformers; wei2023chainofthoughtpromptingelicitsreasoning), which limits reasoning diversity and restricts exploration of multiple valid solutions (naik2024diversitythoughtimprovesreasoning; yu2024flow).
Diffusion models (ho2020denoising), originally introduced for generation in continuous domains like images, have recently gained attention in text generation for their ability to maintain global coherence and enable iterative refinement (ye2024diffusionthoughtschainofthoughtreasoning; nie2025largelanguagediffusionmodels; lou2023discrete; yu2025dimple; weligalle2025discrete; sahoo2024simple; gulrajani2023likelihood). Moreover, prior works have explored continuous or latent diffusion for language generation (li2022diffusion; lovelace2024diffusion; zhang2023planner; lovelacestop; cetin2025large), operating diffusion in latent spaces obtained from text autoencoders or token-embedding spaces. Existing works largely emphasize the parallelization properties of diffusion models (israel2025acceleratingdiffusionllmsadaptive; nie2025largelanguagediffusionmodels; weligalle2025discrete) or evaluate fluency in text generation (li2022diffusion; lovelace2024diffusion; zhang2023planner; lovelacestop). Arguably, a more important direction is to ask: How can these approaches enhance the reasoning capabilities of LLMs? We focus on one particularly promising capability: the ability to self-correct and refine reasoning chains at semantic levels in latent space. As shown in Fig. 1 (top right), such self-refinement cannot be achieved by discrete diffusion language models that merely transit into masked tokens.
@@
To address this limitation, we introduce LaDiR (Latent Diffusion Reasoner), a flexible reasoning framework that encodes high-level semantic representations of reasoning steps into continuous latent tokens via a Variational Autoencoder (VAE) as latent thought tokens, and trains a latent diffusion model over them to perform reasoning. This bridges the gap between surface-level token refinement and deeper semantic reasoning. After the reasoning process, the model generates final answer tokens conditioned on the generated latent thought tokens. Unlike prior latent diffusion works for text generation (li2022diffusion; lovelace2024diffusion; zhang2023planner), which focus on fluent text generation, our framework is explicitly designed for latent reasoning: it learns causal dependencies across reasoning steps through blockwise diffusion, propagates answer correctness signals back to latent tokens.
Our proposed paradigm establishes a new reasoning framework as a post-training method, bringing several distinctive advantages. First, the iterative refinement ability of diffusion enables a better trade-off between accuracy and test-time compute, as additional denoising steps can be flexibly allocated to improve performance. Second, our framework introduces a diversity-guidance mechanism that applies repulsive forces during diffusion inference, pushing latent trajectories apart within a batch to explore multiple diverse reasoning paths, where as AR models tend to collapse to similar trajectories. Finally, leveraging a VAE-based latent space enhances interpretability over continuous diffusion models, making the reasoning process more transparent and readable.
Experimentally, we demonstrate that diffusion-based latent reasoning is not only more accurate but also qualitatively different from prior approaches. On math reasoning benchmarks, including GSM8K (cobbe2021training) and MATH (hendrycks2021measuring), where Coconut (hao2024traininglargelanguagemodels) fails to surpass AR CoT supervised finetuning (SFT), LaDiR consistently outperforms it on average across 7 benchmarks with the LLaMA 3.1 8B model (dubey2024llama). This suggests that modeling reasoning at the semantic level, rather than at the token level, may lead to more faithful intermediate steps that accumulate into stronger final answers. Moreover, on the Countdown planning task, LaDiR shows over 30% absolute improvement in both Pass@1 and Pass@100, indicating that latent thought tokens potentially enhance global planning ability, while parallel diversity exploration enables the model to generate diverse reasoning paths. Together, these findings suggest that diffusion-based latent reasoning provides a principled way to balance accuracy and diversity—key ingredients for advancing beyond sequential autoregressive reasoning.
@@ 2 Preliminaries
This section introduces key concepts and notations in VAE and latent diffusion models (rombach2022high). Detailed formulations and background information are provided in Appendix B.
@@ 2.1 Variational Autoencoder
A Variational Autoencoder (VAE) (kingma2013auto) learns a latent representation of data by balancing reconstruction accuracy and prior regularization. Let $x\in\mathbb{R}^{L\times d_{x}}$ denote a sequence of token embeddings with length $L$ and embedding dimension $d_{x}$, and $z\in\mathbb{R}^{M\times d_{z}}$ denote latent representations with $M$ latent tokens of dimension $d_{z}$. We adopt the $\beta$-VAE (higgins2017beta), where a scaling factor $\beta$ controls this trade-off:
@@
Here, $q_{\phi}(z|x)$ is the encoder distribution parameterized by $\phi$, $p_{\theta}(x|z)$ is the decoder likelihood parameterized by $\theta$, and $p(z)=\mathcal{N}(0,I)$ is the prior distribution over latents.
@@ 2.2 Latent Diffusion and Flow Matching
Latent diffusion models (ho2020denoising; rombach2022high) generate data by denoising latent variables from Gaussian noise in the latent space of a VAE, which preserves high-level semantic structure. Diffusion can also be viewed as a continuous-time generative flow trained via flow matching (lipman2022flow), which we adopt as our primary framework for its superior performance (see Appendix C). The training and inference processes are as follows:
@@ 2.3 Block Diffusion
To support flexible and variable-length sequence generation, we employ a block diffusion scheme (arriola2025block) that integrates autoregressive modeling with diffusion. Instead of applying diffusion to individual latent tokens or full sequence, the sequence is divided into contiguous blocks, and diffusion is performed at the block level. This hybrid design retains the open-ended generation of autoregressive models while introducing global coherence within each block. See Appendix B.4 for details.
@@ 3.3 Reasoning Model Training
While $f_{\psi}$ learns to predict latent reasoning trajectories, to avoid explicitly decoding these steps through the VAE decoder for efficiency during inference, we use the same transformer backbone $\psi$ with a LM head to autoregressively predict answer text tokens conditioned on the latent reasoning blocks. To this end, given the question $q$, the reasoning blocks $\mathbf{Z}^{(\leq B)}$, and the past answer tokens $y_{<w}$, the model predicts the next answer token $y_{w}$ with distribution $p_{\psi}(y_{w}\mid q,\mathbf{Z}^{(\leq B)},y_{<w}).$ The training objective for those answer tokens is the cross-entropy loss:
@@
To explicitly control the number of latent blocks, we introduce a special binary classification head on top of the same LLM transformer backbone $\psi$. It predicts whether the next block begins with a <SOA> (start-of-answer) or <BOT> (begin-of-thought) token whenever an <EOT> (end-of-thought) token is generated. Formally, let $\tau$ index positions of <EOT> tokens in the output. For each $\tau$, the model produces a distribution $p_{\psi}(s_{\tau}\mid q,\mathbf{Z}^{(\leq B)},y_{\leq\tau}),s_{\tau}\in\{\texttt{<SOA>},\texttt{<BOT>}\},$ and we minimize the corresponding classification loss, which supervises the model to predict special tokens given the question $q$ and latent reasoning blocks up to position $\tau$:
$\mathcal{L}_{\mathrm{Spec}}=-\sum_{\tau\in\mathcal{T}_{\text{EOT}}}\log p_{\psi}(s_{\tau}\mid q,\mathbf{Z}^{(\leq B)}).$ (4)
@@ 3.3.2 Stage 2: Rollout training
After Stage 1, there is a mismatch between training and inference. During inference, the model must be conditioned on previous self-generated latents without access to oracle latents, suffering from error accumulation issue. To address this issue, Stage 2 adopts an rollout training. We keep the same number of blocks $B$ as in the ground truth, but instead of conditioning on oracle latents, the model generates its own latents $\tilde{\mathbf{Z}}^{(1:B)}$ from random noise using a fewer denoising steps (i.e., $50\rightarrow 10$, following FlowGRPO (liu2025flow)). We keep the gradients on $\tilde{\mathbf{Z}}^{(1:B)}$ during denoising, allowing answer supervision to backpropagate through the trajectory and directly shape latent predictions. To avoid latent collapse as in Coconut w/o curriculum learning (hao2024traininglargelanguagemodels), we keep the flow matching loss. Therefore, the training objective is same as Eq. 5.
@@ 3.4 Reasoning Model Inference
    Finally, the diversity-guided prediction combines the base model output with the repulsion gradient, in a form analogous to classifier-free guidance (ho2022classifier): $\hat{z}_{t-1}\;=\;f_{\psi}(\mathbf{x}_{t},t,x)\;+\;\gamma_{t}\,\mathbf{F}(z)$, where $f_{\psi}(\mathbf{x}_{t},t,x)$ is the model’s prediction at step $t$.
@@ 4.1 Mathematical Reasoning
We fine-tune pretrained LLMs on the DART-MATH dataset (tong2024dart), a large-scale dataset synthesized to enhance mathematical reasoning. For evaluation, we adopt two in-domain benchmarks, Math (hendrycks2021measuring) and GSM8K (cobbe2021training), and five out-of-domain benchmarks to assess generalization: College-Math (tang2024mathscale), DeepMind-Math (saxton2019analysing), OlympiaBench-Math (he2024olympiadbench), TheoremQA (chen2023theoremqa), and Fresh-Gaokao-Math-2023 (tang2024mathscale). Detailed dataset descriptions are provided in Appendix E.
@@
We compare our approach against a diverse set of reasoning methods spanning both autoregressive and diffusion-based paradigms. For autoregressive models, we include Sol-Only, trained only on question–solution pairs without intermediate reasoning steps, and CoT, trained with full chain-of-thought supervision. We further evaluate several latent reasoning methods: Implicit CoT (iCoT) (deng2023implicit), which gradually removes explicit CoT tokens through curriculum learning; Pause Token (goyal2023think), which introduces a learnable pause token to provide additional computation before answering; Coconut (hao2024traininglargelanguagemodels), which leverages hidden states as latent reasoning tokens through curriculum learning; and Discrete Latent (su2025token), which compresses a span of text (e.g., 16 tokens) into discrete latent codes via VQ-VAE for reasoning. For diffusion-based language models, we compare with the open-sourced LLaDA 8B (nie2025largelanguagediffusionmodels), evaluated both with and without SFT. To directly compare with prior latent diffusion methods for language generation, we also evaluate LD4LG (lovelace2024diffusion) and PLANNER (zhang2023planner). For fair comparison, we use FLAN-T5 as the encoder and LLaMA-3.1-8B as the decoder, and train them on the same reasoning datasets. We utilize LLaMA-3.1 8B (dubey2024llama) as the backbone model for our framework as well as for the AR baselines for fair comparison.
@@
Masked Diffusion Methods - LLaDA 8B
@@ Results
As shown in Table 1, our method achieves the strongest overall performance on average, improving the average pass@1 accuracy by 2% over the best prior latent approach. Also, LaDiR achieves higher Pass@100 across all benchmarks, with a 6.1% absolute gain over AR CoT SFT on average (see Table 8 in Appendix). Compared to text-based CoT baselines, our latent reasoning consistently yields more robust solutions, particularly on harder benchmarks such as DM-Math and College-level datasets, where direct text reasoning often struggles with long-horizon consistency. This suggests that reasoning in a latent space at semantic level learns more abstract reasoning patterns. Compared to prior latent approaches (e.g., Coconut), the latent diffusion objective provides a more principled objective for modeling continuous trajectories, leading to stronger generalization to out-of-domain settings such as TheoremQA. Also, incorporating the stage 2 rollout training notably improves performance across all benchmarks, showing its effectiveness in mitigating error accumulation. Moreover, prior latent diffusion methods of LD4LG and PLANNER perform poorly on reasoning tasks, indicating that effective latent reasoning requires more than architectural changes, such as blockwise variable-length diffusion and rollout training for answer alignment. Taken together, these results indicate that LaDiR combines the interpretability benefits of CoT-style reasoning (see Appendix D) and expressiveness of continuous latent space, producing generalizable reasoning traces.